OpenStudy (anonymous):

How do I do this? 2y^1/3∙ 6y^1/3

OpenStudy (anonymous):

add the exponents

OpenStudy (anonymous):

\[\large 2y^{1/3} \times 6y^{1/3}\]

OpenStudy (anonymous):

also \(2\times 6=12\) so \[12y^{\frac{1}{3}+\frac{1}{3}}\]

OpenStudy (anonymous):

So I would have \[12y ^{2/3}\]

OpenStudy (anonymous):

yes